The absorbance (A) of the ibuprofen solution is concentration-dependent (c, µg / ml) as follows:
A = 0.002391 · c - 0.003.
a) What is the concentration of the solution when the absorbance is 0.735?
b) What is the absorbance at 0.593 mg / ml solution?
Answer #1

A=0.002391.c -0.003 A = 0.735 0.735.= 0.00239jcao.003 0.002391 € = 0.735 + womes 0.003 c= 0.738/0.002391 c = 308.66 ng/ml b)